First Release

TransAmiga

TAFileTagger Documentation

Jürgen A.Lamers December 3, 1995

TAFileTagger  2  Jürgen A.Lamers

Contents

1. Preamble 3 2. Command Overview 3

2.1 Hints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3. Use 6

3.1 Prerequisites . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Explanatory remarks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 Utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4. Thanksgiving 12 5. Warranty 12 6. Future 12 7. Contents of the Archiv 13 8. Registration 15 9. Address 15 10. Technical Hints 16 11. Known Bugs 16

TAFileTagger  3  Jürgen A.Lamers

1. Preamble

Tallyho TransAmiga-SysOp/SysOma ! At first I would like to give a short overview to this program. It bids for the user a comfortable replacement for the normal filemenu :

• Download,

• Upload,

• Private Userarea without changes to the File.cfg,

• Sysop Filearea without changes to the File.cfg,

• Online FRequest Handling,

• Sort the Files.bbs online,

• Viewing texts,

• Viewing archivs, therein reading texts,

• Easy change to another area,

• User can ”’exclude”’ an area,

• Language sensitive,

• Groupstructure. The TAFileTagger is integrated in the TransAmiga environment as far as possible. So every

filemovement of the user is registered on his/her account. I will explain the extras by using and installing this program at the individual commands.

In contrast with the registered version this version has no functional restrictions.

2. Command Overview

• Scroll-Commands

• Area-Commands

• File-Commands

• Common Commands

• Sysop-Commands Scroll-Commands

[figure: 1; Ende; Bottom of the list]
[figure: 7; Pos1; Top of the list]
[figure: 2; ↓; A line down]
[figure: 8; ↑; A line up]
[figure: 3; Bild^{↓}; A side down]
[figure: 9; Bild^{↑}; A side up]
[figure: J]+[ [figure: a].. [figure: z]]  Alphabetic jump
[figure: J]+[ [figure: 0]...]  Directly choice of the area at the group- and areamenu
[figure: .]  Alias to [figure: J]

Area-Commands

[figure: A; -; Change to the areamenu]
[figure: G; +; Change to the groupmenu]
[figure: <; 4; ←; Previous area]
[figure: >; 6; →; Next area]
[figure: ↩]  Choice of group/area
[figure: Q; 5; Terminate]
[figure: P]  Jump to the private filearea
[figure: \]  Jump to the sysop filearea

File-Commands

[figure: M] Mark/Unmark the file below the cursor
[figure: C] Complete abrogate marking
[figure: D] Download the marked files
[figure: X] Upload a FileRequest File
[figure: K] Crunch and download the marked files
[figure: U] Upload
[figure: N] Give a new description to your upload
[figure: I] Information to given file, long description e.g.

[figure: R]

Take a view at a textfile or an archivfile. In an archiv is the same handling as in the main program. Except for ARJ-Files, you can read textfile while you are seeing an archiv-listing.

Common Commands

[figure: -]  Everytime one screen back until the end.
[figure: S]  Get the sortmenu.
[figure: O]  Shows your sort options.
[figure: P]  Get the fileprotocolmenu.
[figure: Y]  Filestatus of the user.

[figure: ?; H; Help; Aid.]

Sysop-Commands If you have the right level, you can switch on the Sysop-Commands with [figure: \]. At the moment the following commands are implemented:

[figure: E] Erase the file beneath the cursor.
[figure: N] Give a new description to the file.

2.1 Hints

Most of the commands are shown in the bottom line and all commands related to the actual menu are explained on the Helpscreen. By the way : All visible commands are executable.

Mark and abrogate the marking is only possible in TAFileTagger, there is no crossreference possible to other choices for download or filelist.

View File :

Because of the same handling through all functions of the program, the management of the filetype-recognition is selfprogrammed. The following archiver are supported :

LHA, LZX, ARJ, ZIP, ZOO, ARC.

Thereby the cruncher had to be in the path, I do not take the entries from file.cfg cause of an ARexx-Entry normaly to this commands.

If you have problems with the archiv-handling, give me a short hint. With the cruncher I have on my machine, it operates faultless.

Upload :

Before the upload an indexscan is possible: the user is asked after the files he/she wants to upload. The named file will be searched in the filebase to check, whether he/she does not take an existing file to your BBS.

Is there any file found with the same name, thus appears a prompt with the entry of the file from files.bbs inclusive which area. The user gets a short question, but none effect is shown.

After some upload, I will ask the user to give a description for his/her upload. There is no limit to the description! This description is also saved in the *.desc-File.

SysopArea:

Start the description at first with ’/’ I will move this file to the given Sysopdirectory. The default directory is ’BBS:SysopFiles’, which I create, if it is not present. At the same way to the private file area you do not have to change the File.cfg.

UploadDataFile:

After every upload, I create a file with the following contents:

<AreaName>,<UserName>,<FileName>,<Size>,<Date>,<Desc>

<Date> = DD MMM YY

You will find this file in the BBS-System directory with the name UserUploads.dat?. The asterix gives you the counting of the existing UploadDataFiles.

Download:

Here you get the possibility of using a built-in special ’protocol’, this shortens the long filenames, which are possible at the Amiga, at 8.3.

Sort:

If anyone wants to resort the files.bbs online, he/she can do this with the sortmenu. It is possible to sort the list by filename (default), date, size and comments (all reversed or natural).

Take one of these sortings it will be the default until ending TAFileTagger. These sort options are saved as <UserName>.sort.<line> and <UserName>.drift.<line> in ENV:.

Magictextfiles:

I support the magictextfile PreUpload, PreDownload and take myself additional textfiles:

PreCheckIndex You can give additional information to the CheckIndex-Questionary
PostUpload  Thanks for his Upload
Arg  I will send this file, if the user wants to read a binary file
PreMarkLimit Hint, that the User can not mark more files.
PreTimeLimit Hint, that download all his marked files will overdraw his online time.
PreRatio  Hint, that he is at the limit of his Upload:Download-Ratio

MagicRexxScripts:

Supporting to the ARexx-Scripts Upload.trans and Download.trans is a little bit difficult. You must add this command to these scripts : ’ADDRESS ’TRANSAMIGA.<line>” and copy them to Upload.tagger.<line> and Download.tagger.<line>

Help-Screens:

During every HelpScreen the user is able to download a global helpfile named Tagger.hlp in your bulletin directory.

TaggerMainHelp  The separated Helpfiles to the Menus.

TaggerListHelp TaggerArcHelp TaggerAreaHelp TaggerPrivateHelp TaggerFreqHelp You can edit and name these Helpfiles to your given userlevels and languages.

3. Use

3.1 Prerequisites

Programs  : Kickstart 2.0 and above, TransAmiga 1.12 and above

Directories : The program ascertains the systemdirectories by TransAmiga on his

own way, additionaly T: and ENV: are used.

Other Things : Nerves.

The TAFileTagger should be installed in the main menu. Versions older than TransAmiga 1.11 (final release) alight the user in the actual filearea, although they choose Quit.

By TransAmiga 1.11 you can use TAFileTagger also as a complete replacement for the filemenu, instead of renaming the script TAFileTagger.trans in File.trans.

3.2 Explanatory remarks

Call :

TAFileTagger has following parameters :

LINE  <n>
CONFIG  <String>

FIRSTAID GROUP

SPECIALACCESS  <n>
GUEST  <n>
PROTOCOL  <n>
UPLOADDIR  <String>

CHECK UPTIME UNAME

SYSOPDIR  <String>

UPDATAFILE NOTIFY

TAG  <String>
ADDR  <n>
PRIVDIR  <String>
PRIVACC  <n>
FREQCFG  <String>
FREQACC  <n>

RATIO NOREAD

SETUPDIR  <String>

<n> : Value (s. TAFileTagger.trans)

LINE <n>

With which TransAmiga-Line shall I operate, this parameter is mandatory.

CONFIG <String>

Argument to give the configfile. You are able to use one configfile by two or more lines. The following way is implemented to handle the configfile:

1. reading BBS:Configs/TAFileTagger.cfg

2. parsing the CLI-parameters

a) If there is any config-parameter, TAFileTagger read this file. Pay attention: CLI-Parameters before these are overwritten. And on the other side parameters from the configfile are overwritten by the following CLI-Parameters!

FIRSTAID

Toggle to get a helpscreen at the beginning. It should be neccesary in the first weeks and to FirstCallern :-).

PROTOCOL <n>

Number of the filetransferprotocol from File.cfg to shorten filenames at 8.3. You should add an entry in the File.cfg like :

      ZModem (PC-Cut)
      1
      BBS:Bin/TrXPR %n zmodem T?,OY,B16,AN,DN,SN s %l
      BBS:Bin/TrXPR %n zmodem T?,OY,B16,AN,DN,KN,SN,RN,P%f r ram:

NOREAD

Disabled the feature to read a textfile or an archiv, if the userlevel is lower than the download access in the actual area.

PRIVDIR <String>

The main directory to all the private directories of the users. If these directories do not exists the TAFileTagger will create them. You can not use spaces in the name of these directories, because TrXPR is not able to handle parameter with quote.

PRIVACC <n>

Userlevel of access to the private filearea.

FREQCFG <String>

Configfile to enable magics during online filerequest. I take the same syntax as FFRS, so you can give me the FFRS-Configfile. If you do not know FFRS, you can create a file with the following syntax:

MAGIC FILE <magicname> <filename with path> [<password>]

FREQACC <n>

Userlevel of access to the online filerequest command.

UPLOADDIR <n>

Directory in which all uploads are placed. The program updates the files.bbs, creates *.desc, the description in the files.bbs gets an additional entry ’Upload: ¡UserName¿’.

UNAME Disable the additional entry ’Upload: <UserName>’.

UPTIME

Switch to tell TAFileTagger to save the upload time of the user (Default: OFF). Well, if MAXTIME is lower than 255 this option has no effect.

CHECK

Toggle to switch on an indexscan before upload. Before using this command you must create an indexfile with TAFileIndex. After each alteration in the fileareas you had to update this indexfile!

If there was any upload during TAFileTagger, it sets an Environment-Variable ’TAUpload’, which you can check in the LogOff.trans.

SYSOPDIR Path to the sysop file area (Default: BBS:SysopFiles).

SETUPDIR Path to the UserSetup directory for AreaExclude (Default: BBS:System/Setups).

UPDATAFILE Switch to disable the datafile after an upload.

NOTIFY

Switch to send a message to the Sysop, if someone has taken an upload. (Default: OFF)

TAG <String>

FidoTag, if you do not want to find the Notify-Message in your Netmail Folder (e.g. WUIVENDRIET.NEWFILES).

ADDR <n> That is the counting number of your address from Message.cfg (Default : 1).

GROUP

Toggle to switch on the groupstructure.

Structure of the Groups:

If you want to add a groupstructure above the areastructure by TransAmiga, thus you must take the following changes respectively additions in the File.cfg:

The groups are defined by the areaflags, this means all areas in one group have the same areaflag. The best way is to list these areas in the individual groups linear. The groupname has to be in the first area of the group, using pad1-entry of the normal structure. In the same line, separated with #, you can set a description to the group, this description is shown in the groupmenu. Pad2 is also used as a description for the area.

Example :

************************************************************************
Wuivend Riet Information                : Areaname
FILES:WuivendRiet/                      : path
0                                       : DownloadAccess
0                                       : UploadAccess
0                                       : AreaAccess
0                                       : AreaFlags, now Groupnumber
Infos zur Mailbox # Wegweiser           : title and comment for this group
Wichtige Sachen fuer Neue User          : comment for this Area
Hilfe fuer DFUe-Anfaenger
FILES:Greenhorn/
0
0
0
1
DFUe-Programme # Tools und Texte
Texte fuer den Einstieg
DFUe Utilities
FILES:Applications/
0
0
0
1
                                        : now a blank pad1
Packer, Terminal-Programme
C't Listings
FILES:ct/
5
10
0

2
Service und Sourcen                     : comment is not necessary
Listings und Inhaltsverzeichnisse
Official Media Agency
FILES:OMA/
0
255
0
2
                                        : Areacomment also
************************************************************************

Making accesible a group only a certain userlevel, you must enter in the first area of this group the wished level in the field areaaccess, so the complete group is not shown for the user with a lower level.

Remember, that new users respectively guests receive the analogs flags for the groups.

SPECIALACCESS <n>

With this parameter, I give guests the possibility to download in some areas (Upload is not possible as guest). These Areas are labelled by Download-Access=0. I think that your care of the SPECIALACCESS is smaller than the the registered users. After leaving TAFileTagger the userlevel resets to 0.

GUEST <n>

Compatible with my TAGuest-Program. Level with the same affect as a FirstCaller.

3.3 Utilities

TAFileIndex Create an indexfile for TAFileTagger. This file is used if you want to prevent unneccessary uploads from your users (like old versions)

Call: TAFileIndex <TransAmiga configfile>

(DEFAULT: bbs:configs/Transmiga.cfg)

TaggerCfgWrite WB-Application to modify the TAFileTagger-Configfile

Call: TaggerCfgWrite <TAFileTagger configfile>

(DEFAULT: bbs:configs/TAFileTagger.cfg)

TaggerConCfg CLI-Application to modify the TAFileTagger-Configfile, so you can take changes within TrShell.

Call: TaggerConCfg <TAFileTagger configfile>

(DEFAULT: bbs:configs/TAFileTagger.cfg)

FileCfg2Txt Extract only the path from File.cfg.

Call: FileCfg2Txt CfgDir/K,PreTxt/K,PostTxt/K

CfgDir: (DEFAULT: bbs:configs)
PreTxt: Text before the path
PostTxt: Text after the path

Example:

FileCfgTxt PreTxt DIRECTORY PostTxt ”’ FREE”’

TAAreaExclude Onlineprogramm within the user can exclude not wished fileareas.

Call: TAAreaExclude Line/A/K/N,SetupDir/K
Line:  BBSLine
SetupDir: Directory where to store the Userdatas, must be the

same as you give to TAFileTagger. (DEFAULT: bbs:system/setups)

You will find an arexxscript with the same feature in the rexx-directory.

UpdateExcl Run this program, if you have changed the fileareas. But of course I need the old fileconfig to do my job.

Call: UpdateExcl OldCfg/A,NewCfg/A,SetupDir/A,Verbose/S
OldCfg: Old File.cfg
NewCfg: New File.cfg
SetupDir: Directory where to store the Userdatas
Verbose: Verbose-Mode

4. Thanksgiving

I would like to thank Sami Radwan for his Oberon-Source from the TransAmiga-Library and the AMOK-Club for their untiring job in reference to Modula-2 and Oberon with the Amiga.

But also I should not forget Andreas Przystaw and would like to thank him, for his help with patience and knowledge in the Beta-phase.

And a general Thank You to the user of my BBS.

5. Warranty

For harms and other results by using these programs I take NONE guarantee.

This software is subject to the ”’Standard Amiga FD-Software Copyright Note”’ It is SHARE-WARE as defined in paragraph 4s. For more information please read ”’AFD-COPYRIGHT”’ (Version 1 or higher).

6. Future

• Debugging :-(

• Incorporate other proposals.

• In also future integrate an ARexx-interface.

7. Contents of the Archiv

bin/ -

TAFileTagger  - Mainprogram
TAFileIndex  - Utility to generate the indexfile used by Upload-Scan

TaggerCfgWrite - WB-Application to modify the TAFileTagger-

Configfile

TaggerConCfg - CLI-Application to modify the TAFileTagger-

Configfile, so you can take changes within TrShell.

TAAreaExclude  - External Tool to exclude fileareas by the user.

UpdateExcl - Sysop Tool to update the exclude-setups of your

users.

FileCfg2Txt  - Nothing special…

configs/ -

TAFileTagger.cfg  - Example Config-File
Tagger.lang  - Language File (german)

Tagger.english.lang - Tagger.Deutsch.lang - text/ -

TaggerMainHelp.ANSI - Helpfiles.

TaggerListHelp.ANSI - TaggerArcHelp.ANSI - TaggerAreaHelp.ANSI - TaggerPrivateHelp.ANSI - TaggerFreqHelp.ANSI - TaggerSysopHelp.ANSI - rexx/ - TAFileTagger.trans - ARexx-Script to launch TAFileTagger by

TransAmiga

PrivUser.logon - ARexx-Script to generate the private FileArea during

Login.

FExclude.trans  - ARexx-Script, do the same as TAAreaExclude
Download.tagger  - TAFileTagger-MagicRexxScript ala Download.trans.
Upload.tagger  - TAFileTagger-MagicRexxScript ala Upload.trans.

examples/ -

Logon.trans  - Example-Script
LogOff.trans  - Example-Script
Main.rx  - Example Config-File
UpdateTAIndex  - Example Script to update the indexfile

docs/ -

TAFileTagger.doc  - English documentation.
TAFileTagger.dok  - Deutsche Dokumentation.
TAFileTagger E.dvi  - Documentation (en).
TAFileTagger D.dvi  - Documentation (de).
RegForm E.txt  - Plain-ASCII-File for registration (en).
RegForm D.txt  - Plain-ASCII-File for registration (de).
TAScrollerDoor.lha  - The son of TAFileTagger.

No changes are allowed to this archiv !!

8. Registration

Well, some little differences between registered and unregistered version exist, these are shown by using: NodeLocation is set to

’<UserName> is using TAFileTagger[UNREGISTERED]’

as readily appears

’REGISTERED TO [UNREGISTERED]’,

after terminate a small requester appears and the user has to wait 15 seconds.

It would enjoy me, if the user persuades the sysop to use this program in the registered version. The registration fee amounts to DM 25,-.

This fee contains registered versions of: TAFileTagger TAFileScroller - ANSI-FileScroller, it lacks only the option GROUP and runs as re-

placement for filelist.

Updates Hotline

Payment of the registration fee:

Remittance : Jürgen A.Lamers Postgiroamt Köln KtoNr. : 4102 31-507 BLZ : 370 100 50

Eurocheque DM (Please only in DM) exhibit at : Jürgen A.Lamers Altstr. 112 D-52066 Aachen GERMANY

9. Address

OMA-Software Jürgen A.Lamers Altstr. 112 D-52066 Aix la Chapelle Tel.: +49-(0)241-571489 Fax : +49-(0)241-573145 Box : +49-(0)241-573145

Fido  Juergen A.Lamers  2:2452/121.0
E-Mail jaloma@wuivendriet.fido.de 242:3000/3.0

10. Technical Hints

You will find the following file after and during the usage of TAFileTagger:

<username>.tagger The tagged files, separated with a space. (In T:)
UserUploads.dat<n> Data to the last Uploads. (In BBS:System)

And following Environmentvariables:

<username>.bytes.<line> Count of the tagged Bytes.
<username>.files.<line> Count of the tagged Files.
<username>.sort.<line> Used Sortoption.
<username>.drift.<line> Used Sortdrift.
TAUpload  Set to 1 when the user takes an upload.

11. Known Bugs

• NEVER toggle LocalDisplay while TAFileTagger is running.

• Well, it is possible to chat while TAFileTagger is running. But if you leave the chat, the user will find themself at the main menu and you will find out, that TAFileTagger is still running

  :-(.

• Some problems with AsXpr.

• Some problems to identify binaries and PC-textfiles, so sometimes the user reads a binary … • Not a bug, but I think that the options SPECIALACCESS and GUEST are not neccessary,

so I will remove them by the next release.